What is the de Broglie wavelength, in centimeters, of a 12.4 g hummingbird flying at 40 mph
1
Expert's answer
2020-10-23T09:16:15-0400

We can list down the data available in the question, changing all values to appropriate SI units:

  • h=6.626×10-34Js


  • m=12.4×10-3kg


v=1.20×102mph×0.447m/s/mph=53.64m/s


Now, the de Broglie wavelength is:

λ=(6.626×10-34J.s)/(12.4×10-3kg) ×(53.64m/s)=9.96×10-34m


We can change this to centimeters by multiplying by 100:


λ=9.96×10-32cm